A grain silo is shown below:


Grain silo formed by cylinder with radius 6 feet and height 168 feet and a half sphere on the top


What is the volume of grain that could completely fill this silo, rounded to the nearest whole number? Use 22 over 7 for pi.


19,008 ft3

19,461 ft3

6,336 ft3

453 ft3


1
Expert's answer
2020-07-07T21:04:52-0400

V=total volume, R=6ft, H=168ft.

V=V(cylinder)+1/2V(sphere)

V(cylinder)="\pi" *R2*H=(22/7)*36*168=19008ft3

1/2V(sphere)=(1/2)*(4/3)*"\pi" *R3=(2/3)*(22/7)*216=452.571429ft3

V=19008+453=19461ft3